Circular No. 7349

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 B1 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Cen-

ter, reports observations of a presumed Kreutz sungrazing comet discovered by J. D. Shanklin (Comet Section, British Astronomical Association) in SOHO/LASCO C3 data (still outside the field of the C2 instrument). Astrometry (measured by Biesecker and D. Hammer, reduced by B. G. Marsden) and parabolic orbital elements with assumed q = 0.0052 AU are given on MPEC 2000-B22. Biesecker provides apparent magnitudes brightening from V = 7.36 on Jan. 24.571 to 5.67 on Jan. 24.846 UT.

2000 UT  α_{2000}  δ_{2000}
Jan. 24.388  20^{h}52.^{m}2  −22^{o}30′

CATACLYSMIC VARIABLE IN URSA MAJOR M. Uemura and T. Kato, Kyoto University; R. Novak, Nicholas Coper-

nicus Observatory; and L. T. Jensen, VSNET collaboration team, report that a variable star that was discovered by K. Takamizawa, Saku-machi, Nagano, Japan, in outburst (mag 13.0, T-Max 400 film) on 1997 Nov. 9, was recently found to be in outburst for a second time on 2000 Jan. 13.509 by P. Schmeer, Bischmisheim, Germany, at unfiltered CCD mag 14.0; Schmeer had found the star at mag 17.6 on Jan. 8.463. Takamizawa reports that the variable was fainter than mag ∼ 15 on numerous other dates back to 1994 Nov. 15. H. Yamaoka reports the following position for this new variable: α = 10^{h}43^{m}56^{s}.87, δ = +58^{o}07′32′′.5 (equinox 2000.0). CCD photometry on Jan. 17 by Uemura et al. revealed deep (1.3 mag) eclipses and distinct superhumps, indicating that the star belongs to a rare class of deeply eclipsing SU UMa-type dwarf novae. The orbital and superhump periods estimated from the photometric data obtained between Jan. 17 and 20 are 0.0739 and 0.0762 day, respectively; the star maintained mag ∼ 14.4 during this interval. A. Henden obtained a quiescent magnitude V = 18.381 on 1999 June 23, and the star is present in the USNO-A2.0 catalogue. Uemura et al. are currently conducting an intensive observing campaign. The light curve and other related information can be viewed at http://www.kusastro.kyoto-

u.ac.jp/vsnet/DNe/tmzv85.html.

2000 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green